### v1

# MLLM Machine Unlearning via Visual Knowledge Distillation

Abstract
Recently, machine unlearning approaches have been proposed to remove sensitive information from well-trained large models. However, most existing methods are tailored for LLMs, while MLLM-oriented unlearning remains at its early stage. Inspired by recent studies exploring the internal mechanisms of MLLMs, we propose to disentangle the visual and textual knowledge embedded within MLLMs and introduce a dedicated approach to selectively erase target visual knowledge while preserving textual knowledge. Unlike previous unlearning methods that rely on output-level supervision, our approach introduces a Visual Knowledge Distillation (VKD) scheme, which leverages intermediate visual representations within the MLLM as supervision signals. This design substantially enhances both unlearning effectiveness and model utility. Moreover, since our method only fine-tunes the visual components of the MLLM, it offers significant efficiency advantages. Extensive experiments demonstrate that our approach outperforms state-of-the-art unlearning methods in terms of both effectiveness and efficiency. Moreover, we are the first to evaluate the robustness of MLLM unlearning against relearning attacks.
1 Introduction
Recently, Multimodal Large Language Models (MLLMs) have achieved remarkable progress, demonstrating outstanding performance on various multimodal applications such as image captioning and visual question answering (VQA) [liu2023visual, bai2023qwen]. However, growing concerns over data privacy have become a significant barrier to the widespread application of LLMs and MLLMs. This is because training large models typically involves the use of vast amounts of Internet data, which often contain sensitive personal information such as social security numbers or personal photographs. Moreover, numerous studies [pi2024mllm, li2024digger, cohen2024performance] have shown that large models can memorize portions of their training data and reproduce them verbatim in their outputs. This poses a substantial risk of privacy leakage. Therefore, the advent of data privacy and protection regulations—such as the European Union’s GDPR [regulation2018general] and the California Consumer Privacy Act (CCPA) [goldman2020introduction]—now mandates these large model providers to honor data deletion requests from individuals [dang2021right].
The most straightforward approach to protecting privacy is to discard the trained model entirely, remove the individual’s data from the training set, and retrain a new model from scratch. However, retraining is computationally expensive and resource-intensive, particularly for large models. As a result, the field of machine unlearning [garg2020formalizing, ginart2019making, cao2015towards, gupta2021adaptive, sekhari2021remember, brophy2021machine, ullah2021machine] has emerged, focusing on the challenge of efficiently revising a trained model to forget specific data while preserving performance on the retaining dataset. Nowadays, many large-model-oriented machine unlearning methods have been proposed, such as Gradient Ascent (GA) [thudi2022unrollingsgd], its variations [liu2022privateunlearning], and Negative Preference Optimization (NPO) [zhang2024npo], etc. However, most existing approaches primarily focus on unlearning for LLMs, with relatively limited exploration in the context of MLLMs.
In general, unlearning in LLMs is clearly defined—erasing the knowledge of a specific entity from the model. In contrast, defining unlearning for MLLMs is far more complex, as two distinct forms of knowledge are associated with the entity of interest: textual knowledge, stored within the LLM component, and visual knowledge, or visual patterns, linked to the entity’s appearance. For instance, for a well-known individual, the name (e.g., “Robin Williams”) and textual attributes (such as occupation or home address) constitute the textual knowledge, while the person’s facial appearance represents the visual knowledge. Consequently, there remains ambiguity in defining MLLM unlearning: some studies argue that it should involve erasing both textual and visual knowledge [liu2025MANU], whereas others contend that only the visual knowledge should be removed while preserving the textual knowledge [huo2025mmunlearner].
In this paper, we follow [huo2025mmunlearner] in defining the desired MLLM unlearning objective as erasing visual knowledge while preserving textual knowledge, as this formulation offers greater flexibility. The essence of this definition lies in disentangling visual and textual knowledge, thereby allowing selective removal of either component when necessary. Even when the goal is to erase both textual and visual knowledge, this can be achieved by sequentially performing MLLM unlearning followed by LLM unlearning.
An MLLM typically consists of three components: a vision encoder, an LLM backbone, and a projector that bridges the two. Unlike [huo2025mmunlearner], which fine-tunes all three components for unlearning, we propose to fine-tune only the vision encoder and the projector while keeping the LLM backbone frozen. This design is motivated by some studies that explore the internal mechanisms of MLLMs [cohen2024performance, huang2024commonsenseknowledgeeditingbased, yu2024understanding]. These works decompose the VQA process into two stages: the first involves visually recognizing the entity (termed identification), and the second connects the recognized entity to its stored factual knowledge (termed extraction). It has been shown that, in a well-trained MLLM, the vision encoder and projector primarily handle the identification function, whereas the LLM module is responsible for the extraction function. Therefore, to effectively erase visual knowledge while preserving textual knowledge, we propose to keep the LLM module frozen during the unlearning process.
Note that  [huo2025mmunlearner] conducted an ablation study comparing such partial fine-tuning setting (fine-tuning only the vision encoder and projector) with the full fine-tuning setting. Their results demonstrated that the former is inferior to the latter in balancing unlearning effectiveness and model utility. In this paper, we propose a novel scheme, termed Visual Knowledge Distillation (VKD), to harness the advantages of our partial fine-tuning setting.
Specifically, when designing the unlearning objective function, most existing methods focus primarily on the model’s output. Take GA as an example: for the forgetting knowledge, it encourages the model to avoid producing the ground-truth answer, while for the remaining knowledge, it encourages the model to retain correct predictions. We argue that such output-level supervision is well-suited for full fine-tuning settings. However, in the partial fine-tuning scenario, this supervision signal must propagate through the frozen LLM, and thus becomes significantly weakened by the time it reaches the visual module.
To address this issue, we propose to inject the supervision signal directly into the visual module. Concretely, we employ the original MLLM as the teacher model and use its intermediate visual representation (i.e., the output of the projector) as the supervision signal. We then introduce a loss term that encourages the unlearned model (the student model) to align with this signal. We term this approach Visual Knowledge Distillation (VKD), as it distills the teacher model’s visual representations—which predominantly encode visual knowledge—into the student model during unlearning. The empirical results demonstrate that our approach outperforms previous methods in terms of both unlearning effectiveness and model utility.
In addition to its superior unlearning performance, our partial fine-tuning scheme also offers notable advantages in efficiency. Since the knowledge distillation is applied solely to the visual component, the fine-tuning process is highly efficient, requiring updates to only a small portion of parameters. Furthermore, due to its lightweight nature, our approach is well-suited for online unlearning scenarios, where data deletion requests occur frequently and arrive sequentially. In such cases, our method can efficiently fine-tune the visual component on demand for each request.
Beyond evaluating the effectiveness of unlearning, we also examine its robustness. Recent studies have revealed that machine unlearning is vulnerable to the so-called relearning attack, which attempts to recover forgotten data or knowledge using only a few samples from the forgetting set [lynch2024robustunlearning, hu2025joggingmemory]. If the ‘already-unlearned’ knowledge can be rapidly recovered, it suggests that the unlearning process merely obfuscates the knowledge rather than truly removes it. Notably, this is the first work to consider the robustness of unlearning in MLLMs.
We conduct extensive experiments on representative MLLMs, including LLaVA and Qwen-VL, and compare our method with existing MLLM unlearning approaches such as MMUnlearner and MANU. Our approach surpasses these state-of-the-art methods, and interestingly, the proposed VKD module facilitates both the forgetting of target visual knowledge and the retention of non-target visual knowledge. Our method effectively disentangles textual knowledge from visual knowledge. In addition, we evaluate the robustness of our approach against relearning attacks, demonstrating that the forgotten visual knowledge cannot be easily recovered, thereby confirming that our MLLM unlearning method possesses substantial robustness.
2 Related work
Machine Unlearning for LLMs.
Since LLMs may unintentionally memorize their pretraining data, Machine Unlearning (MU) techniques have been introduced to erase privacy-sensitive or copyrighted information from trained models. Gradient Ascent (GA) [thudi2022unrolling] is a representative method that fine-tunes the model to reduce the probability of generating sensitive information. One inherent challenge in unlearning lies in balancing unlearning effectiveness and model utility—that is, erasing the target data while preserving the remaining data. To address this, several GA variants have been proposed to mitigate the impact on remaining data by incorporating KL-divergence minimization [maini2024tofu] or by explicitly reducing the loss on the retained data [rafailov2023direct]. Beyond supervised fine-tuning, several reinforcement learning-based methods, such as Negative Preference Optimization (NPO) [zhang2024negative], have been proposed to further enhance the effectiveness of unlearning.
Machine Unlearning for MLLMs.
Compared to MU for LLMs, research on MU for MLLMs remains in its infancy. SIU [li2024single] explores the erasure of visual patterns in MLLMs using real-world entity datasets through multifaceted fine-tuning. As noted by [yu2024understanding, cohen2024performance], MLLMs exhibit strong cross-modal interactions that entangle visual and textual knowledge, making selective unlearning more challenging and often resulting in substantial unintended knowledge loss. Moreover, even the definition of MLLM unlearning remains controversial: some studies define it as erasing visual knowledge while preserving textual knowledge [huo2025mmunlearner], whereas others argue it should involve forgetting both visual and textual knowledge [liu2025modality]. Nevertheless, existing methods predominantly rely on model outputs as the supervision signal for unlearning. In contrast, our approach leverages intermediate representations through a visual knowledge distillation strategy, providing more direct and effective supervision.
Figure 1: Illustration of our machine unlearning framework for MLLMs. For the target entity, our objective is to (a) forget its visual knowledge while (b) preserving its textual knowledge. For non-target entities, we preserve both (c) the textual knowledge and (d) the visual knowledge through Visual Knowledge Distillation (VKD).
3 Method
3.1 MLLM Unlearning Problem
An MLLM model $M$ typically consists of three main components: a vision encoder $\mathcal{V}$, an LLM module $\mathcal{W}$, and a projector $\mathcal{P}$ that bridges the two. Generally, an MLLM model is trained on a visual instruction dataset $D={(v_{i},q_{i},a_{i})}_{i=1}^{N}$, where $v_{i}$ denotes the input image, $q_{i}$ the textual question, and $a_{i}$ the corresponding answer. Specifically, the training of the model $M$ is carried out through a visual question answering (VQA) task as follows: (1) The image $v_{i}$ is first processed by the vision encoder and the projector, yielding the visual embedding $v^{e}_{i}=\mathcal{P}(\mathcal{V}(v_{i}))$. (2) The question $q_{i}$ is tokenized into a textual embedding $q^{e}_{i}$, which is then concatenated with $v^{e}_{i}$ and fed into the LLM module $\mathcal{W}(q^{e}_{i}\oplus v^{e}_{i})$. (3) The LLM subsequently generates an output that is expected to align with the target answer $a_{i}$.
In the context of MLLM unlearning, we distinguish two types of knowledge associated with the entity of interest: visual knowledge and textual knowledge. We define MLLM unlearning for an entity of interest as the process of erasing its visual knowledge while preserving its textual knowledge. Additionally, both the visual and textual knowledge associated to other entities should remain unaffected.
To evaluate the unlearning performance, we employ both VQA and QA tasks to probe the visual and textual knowledge retained by MLLMs. Specifically, given a data sample of the target entity $(v_{i},q_{i},a_{i})$, we instruct the model to perform both the VQA and QA tasks, where the image $v_{i}$ is omitted in the QA task. If the model fails the VQA task (producing an incorrect answer) but succeeds in the QA task (generating the correct answer $a_{i}$), we regard the model as having forgotten the visual knowledge while preserving the textual knowledge of the target entity. Moreover, for the remaining data samples unrelated to the target entity, the model is expected to successfully perform both the VQA and QA tasks.
Note that conventional machine unlearning methods typically divide the training dataset $D$ into two subsets: the forgetting dataset $D^{f}$, which is associated with the target entity, and the remaining dataset $D^{r}$, which contains all other data. Thus, the unlearning problem can be formulated as:
$\mathop{\min}_{\theta}(-\mathcal{L}^{f}(\theta)+\mathcal{L}^{r}(\theta))$ (1)
where $\mathcal{L}^{f}(\theta)$ denotes maximizing the loss on $D^{f}$, $\mathcal{L}^{r}(\theta)$ represents minimizing the loss on $D^{r}$.
Clearly, this formulation is unsuitable for our MLLM unlearning objective, as the forgetting dataset $D^{f}$ contains both visual and textual knowledge related to the target entity. Therefore, we reformulate the unlearning objective as follows:
$\mathop{\min}_{\theta}(-\mathcal{L}^{f}_{vqa}(\theta)+\mathcal{L}^{f}_{qa}(\theta)+\mathcal{L}^{r}_{vqa}(\theta)+\mathcal{L}^{r}_{qa}(\theta))$ (2)
where $\mathcal{L}^{f}_{\text{vqa}}(\theta)$ denotes maximizing the VQA loss on $D^{f}$, $\mathcal{L}^{f}_{\text{qa}}(\theta)$ represents minimizing the QA loss on $D^{f}$, and $\mathcal{L}^{r}_{\text{vqa}}(\theta)+\mathcal{L}^{r}_{\text{qa}}(\theta)$ represents minimizing the VQA and QA loss on $D^{r}$.
3.2 Our Approach
In this paper, we propose a novel and efficient machine unlearning algorithm to address the MLLM unlearning problem defined in the previous section. Unlike existing unlearning methods that modify the entire MLLM, our approach focuses solely on revising the visual module—that is, the vision encoder and the projector—within the MLLM. This design is motivated not only by efficiency considerations (as adjusting the visual module is significantly more lightweight than updating the entire MLLM) but also by the very definition of the MLLM unlearning objective.
Since modern MLLMs are typically built upon a pretrained LLM backbone, it is widely assumed that textual knowledge is primarily encoded in the LLM, while visual knowledge is largely captured by the visual module [zhu2023minigpt, chen2024internvl]. Moreover, one of our core unlearning goals is to preserve textual knowledge while removing visual knowledge. Therefore, freezing the parameters of the LLM naturally benefits the retention of textual knowledge.
However, when only the visual module is fine-tuned during unlearning, maintaining the delicate balance between unlearning effectiveness and model utility becomes substantially more challenging. As is well known, this trade-off has long been a central problem in machine unlearning. To mitigate it, prior studies have explored strategies such as incorporating KL-divergence regularization to preserve remaining knowledge or applying masking mechanisms to reduce conflicts between forgotten and retained knowledge. We adopt the masking mechanism in our framework; nonetheless, empirical results reveal that it alone is insufficient to meet our requirements. To this end, we introduce a new mechanism—Visual Knowledge Distillation (VKD)—to better balance unlearning and utility preservation.
3.2.1 Visual Knowledge Distillation
It is important to note that all existing unlearning methods generally tackle this balance at the loss level: they maximize the loss on the forgetting dataset $D^{f}$ (to erase target knowledge) while minimizing the loss on the remaining dataset $D^{r}$ (to preserve other knowledge). This paradigm is suitable when the knowledge to be forgotten is stored across the entire model, as the output-level supervision can backpropagate through all model components. However, it becomes less effective in our setting, where the visual knowledge to be erased is concentrated within the visual module, and consequently, only the visual module is fine-tuned during unlearning. Since an MLLM is structured in a cascaded manner—with the visual module feeding into the LLM backbone, the output-level supervision must propagate through the frozen LLM, and thus becomes significantly weakened by the time it reaches the visual module.
In this paper, we propose injecting supervision signals directly into the visual module, since it is the only component being updated during unlearning. We argue that such intermediate-level supervision is more efficient and effective than conventional output-level supervision. Specifically, we use the vanilla MLLM as the teacher model, extracting its intermediate features from the visual module as supervision signals. The target model undergoing unlearning serves as the student model, which aligns its visual features with the teacher’s supervision signals for the visual knowledge to be preserved (i.e., $D^{r}$). This process, illustrated in Fig. 1(d), enables the model to maintain the general visual representation capability of the teacher while selectively unlearning the undesired visual knowledge. Thus, our VKD unlearning approach can be formulated as:
$\mathop{\min}_{\theta}(\mathcal{L}_{output}+\beta\mathcal{L}_{\text{VKD}})$ (3)
[…]
$\mathcal{L}_{output}=-\textbf{m}\odot\alpha\mathcal{L}^{f}_{vqa}+\mathcal{L}^{f}_{qa}+\mathcal{L}^{r}_{vqa}+\mathcal{L}^{r}_{qa}$ (4)
where $f^{r}_{vqa}$ indicates the student model’s visual feature for VQA input data, and $t^{r}_{vqa}$ indicates the teacher model’s visual feature for the same inputs.
Notably, the term $f^{r}_{vqa}-t^{r}_{vqa}||^{2}$ encourages the unlearned model to preserve visual knowledge unrelated to the target entity. Our empirical results indicate that when only the visual module is fine-tuned, the key to balancing unlearning effectiveness and model utility lies in retaining the visual knowledge unrelated to the target entity, while the erasure of target visual knowledge can be sufficiently achieved through output-level supervision.
Our approach involves two weighting parameters. Specifically, $\beta$ is introduced to balance the output-level objective and the VKD objective. The parameter $\alpha$ is employed to trade off among the visual knowledge to be forgotten ($\mathcal{L}^{f}_{vqa}$), and the visual and textual knowledge to be preserved ($\mathcal{L}^{r}_{vqa}$, $\mathcal{L}^{f}_{qa}$ and $\mathcal{L}^{r}_{qa}$).
3.2.2 Selective Forgetting
As shown in Eq 4, we follow the method [huo2025mmunlearner] that leverage a masking scheme as performing unlearning, where m is a boolean mask that selectively updates parameters, and $\odot$ denotes the Hadamard product. Specifically, the mask m is referred to as the weight saliency map, as it quantifies how salient each parameter is with respect to a given dataset $D$.
In addition to saliency-based masking, another selective forgetting technique is neuron pruning, which identifies neurons closely associated with the knowledge to be forgotten and achieves unlearning by pruning them. In this paper, we integrate these two selective forgetting strategies. They can be executed either in parallel or sequentially. Our experiments demonstrate that the best performance is achieved when pruning is applied first, followed by masking-based fine-tuning.
It is worth noting that in our approach, both neuron pruning and weight masking are applied exclusively to the visual module—specifically, the vision encoder $\mathcal{V}$ and the projector $\mathcal{P}$. This is motivated by recent studies on localizing factual memory within Transformer architectures. Prior research [liu2025MANU] has shown that MLP layers serve as primary knowledge storage components in Transformer architectures [geva2021transformerfeedforwardlayerskeyvalue, meng2023locatingeditingfactualassociations], whereas the self-attention modules mainly function as mechanisms for information propagation [elhage2021mathematical].
In particular, for vision encoders such as the CLIP model, it has been observed that the high-dimensional inner projections within the feed-forward layers play a crucial role in storing visual knowledge, whereas the representations derived from the self-attention layers do not [ghiasi2022visiontransformerslearnvisual]. Furthermore, it has been also observed that earlier layers learn textural attributes, whereas deeper layers learn high level object features or abstract concepts [ghiasi2022visiontransformerslearnvisual]. Therefore, in our approach, both neuron pruning and weight masking are applied exclusively to the vision encoder’s MLP layers, with particular emphasis on the deeper layers.
4 Evaluation
4.1 Experimental Setup
To accurately assess whether the target knowledge has been successfully unlearned from the MLLM, we typically do not perform unlearning directly on the original pretrained model (e.g., LLaVA-1.5-7B), as it is uncertain whether the model initially possesses the target knowledge. In practice, most unlearning benchmarks adopt a two-stage procedure: first, a fictitious dataset representing the target knowledge (e.g.,, synthetic personal profiles) is curated, and the original MLLM is fine-tuned on this dataset to obtain a vanilla model that explicitly contains the target knowledge. Then, the unlearning algorithm is applied to this vanilla model, and the resulting unlearned model is evaluated to assess the effectiveness of the unlearning process.
Methods MLLMU-Bench CLEAR
Forget VQA Forget QA Retain VQA Retain QA Realworld VQA Realworld QA Forget VQA Forget QA Retain VQA Retain QA Realworld VQA Realworld QA
Acc($\downarrow$) Acc($\uparrow$) Acc($\uparrow$) Acc($\uparrow$) Acc($\uparrow$) Acc($\uparrow$) Acc($\downarrow$) R-L($\uparrow$) Acc($\uparrow$) R-L($\uparrow$) Acc($\uparrow$) Acc($\uparrow$)
LLaVA-1.5-7B
Vanilla 45.8 38.4 45.2 37.5 47.4 54.9 63.3 0.367 54.0 0.352 53.7 85.4
GA 43.2 32.5 45.0 32.2 47.0 55.0 57.4 0.153 48.4 0.176 51.8 83.4
GA_Diff 40.0 33.6 44.3 31.5 46.6 53.6 47.3 0.197 43.4 0.220 47.7 73.5
KL_Min 42.4 33.6 44.9 32.0 47.4 54.6 40.4 0.270 38.1 0.274 51.5 82.8
NPO 43.2 33.6 44.9 32.2 47.0 55.0 40.4 0.285 38.6 0.282 52.9 83.4
MANU 41.6 33.6 43.2 32.1 46.3 54.8 39.8 0.263 25.8 0.224 49.8 68.3
MMU 31.2 34.2 44.2 35.1 46.7 54.9 36.2 0.348 46.6 0.338 52.3 84.1
Ours 29.8 35.8 45.0 36.1 46.9 54.9 32.7 0.327 49.6 0.346 52.3 84.8
Qwen2-VL-7B
Vanilla 55.2 55.0 56.0 58.6 77.3 77.5 67.0 0.116 70.9 0.098 69.2 91.4
GA 50.4 46.7 51.5 57.6 74.4 77.8 55.3 0.123 62.4 0.083 65.9 86.8
GA_Diff 54.4 52.8 38.8 54.4 74.5 77.0 63.3 0.125 71.4 0.088 69.0 92.7
KL_Min 45.6 45.3 35.9 55.6 74.8 77.1 67.0 0.120 70.9 0.098 68.4 90.7
NPO 49.6 50.4 49.5 53.3 75.2 78.3 62.8 0.103 68.3 0.091 68.9 88.7
MANU 43.2 47.2 43.2 49.4 73.7 76.2 61.3 0.107 65.2 0.094 66.2 87.3
MMU 44.0 54.4 56.0 55.7 75.3 77.3 50.0 0.123 70.9 0.100 68.9 94.7
Ours 42.8 54.6 56.0 56.4 74.9 77.4 48.7 0.125 71.5 0.094 69.0 90.8
Table 1: Overall results of baselines and our approach on MLLMU-Bench and CLEAR benchmarks. $\downarrow$ Lower is better for Forget VQA accuracy (reflecting better forgetting of target visual knowledge), $\uparrow$ Higher is better for the others (reflecting better retention of non-target visual knowledge and textual knowledge). For MLLMU-Bench, the Forget/Retain QA tasks are evaluated using the Acc metric, while for CLEAR, these tasks are evaluated using the ROUGE-L metric.
4.1.1 Datasets and Metrics
Datasets: We employ two benchmark datasets to evaluate unlearning performance in MLLMs. MLLMU-Bench comprises fictitious personal profiles, each paired with a portrait and 18 corresponding questions (i.e., 9 VQA questions and 9 textual QA questions) with multiple-choice options. The dataset is divided into three subsets: Forget, Retain, and Real-world. The Forget Set contains a subset of fictitious profiles designated for evaluating unlearning effectiveness. The Retain Set includes fictitious profiles excluded from the Forget Set, representing the knowledge the model should preserve. The Real-world Set consists of real celebrity profiles distinct from the fictitious ones and is used to assess unlearning entanglement with neighboring real-world concepts, thereby reflecting the model’s utility. CLEAR is built on top of an LLM unlearning dataset, i.e., the TOFU dataset. For each author in TOFU, CLEAR adds several face images and corresponding textual descriptions generated by GPT-4o [achiam2023gpt]. Similar to MLLMU-Bench, CLEAR also comprises Forget, Retain, and Real-world sets.
Metrics: We evaluate the forgetting and retention of textual knowledge using the Question Answering (QA) task, and assess the forgetting and retention of visual knowledge using the Visual Question Answering (VQA) task. The performance on multiple-choice questions is measured by average accuracy, while text generation performance is evaluated using the ROUGE-L metric [lin2004rouge].
Beyond evaluating the effectiveness of machine unlearning, we further investigate its robustness. Prior studies have demonstrated that LLM unlearning is susceptible to relearning attacks, which aim to restore forgotten knowledge using only a few samples from the forgetting set. To the best of our knowledge, our work is the first to explore the robustness of unlearning in MLLMs. Specifically, after performing MLLM unlearning, we fine-tune the unlearned model with a small subset of the forgotten visual knowledge and subsequently assess the extent to which this forgotten knowledge can be recovered through a relearning process.
4.1.2 Models and Baselines
MLLMs: We conduct experiments using open-source MLLMs as base models. Specifically, we use LLaVA-1.5-7B-hf and Qwen2-VL-7B-Instruct, which serve as the foundation for our unlearning experiments. The vanilla models are trained following the official implementations provided by MLLMU-Bench and CLEAR. For each vanilla model, we perform selective unlearning of visual knowledge according to our proposed framework, while freezing the LLM weights.
Baselines: We compare our method with six baseline approaches: GA [thudi2022unrolling], which applies gradient ascent on the forget VQA set $D_{f}$; GA_Diff [liu2022continual], an improved GA variant with a joint loss to balance the forget set $D_{f}$ and retain set $D_{r}$; KL_Min [maini2024tofu], which aligns predictions on $D_{r}$ with the original model while maximizing KL divergence with $D_{f}$; NPO [zhang2024negative], which treats $D_{f}$ as negative-preference data and integrates unlearning into a preference optimization framework with an oracle model fine-tuned on $D_{r}$; MANU [liu2025modality], which introduces modality-aware neuron unlearning by selectively pruning neurons that contribute most to the forget set, preserving model performance while achieving balanced forgetting across modalities; MMUnlearner [huo2025mmunlearner], which uses saliency-weighted updates to selectively protect non-target parameters while forgetting visual knowledge.
All implementations are based on official MLLMU-Bench and CLEAR code, using the same learning rate, forgetting rounds, batch size, and full-precision training on four NVIDIA 4090 GPUs.
Figure 2: The trade-off between the output-level objective and the VKD objective with respect to the VKD strength $\beta$.
Figure 3: The trade-off between knowledge forgetting and preservation with respect to the weighting parameter $\alpha$.
4.2 Main Results and Analysis
4.2.1 Effectiveness of Unlearning
On MLLMU-Bench, our method achieves the lowest accuracy on the Forget VQA task for both LLaVA-7B and Qwen2-VL, surpassing the state-of-the-art method MMUnlearner by 1.4% and 1.2%, respectively. Similarly, on CLEAR, our approach consistently surpasses MMUnlearner by 3.5% and 1.3%, respectively. This illustrates that our approach can effectively erase the visual knowledge associated with the target entity.
Our method also achieves state-of-the-art performance on both the Retain QA task and the Forget QA task, particularly in MLLMU-Bench dataset. This demonstrates that our approach effectively preserves the textual knowledge associated with both target and non-target entities.
Compared to textual knowledge forgetting, retaining the visual knowledge associated with non-target entities remains the most challenging issue. From the results on the Retain VQA and Real-world VQA, our approach shows substantial improvement, outperforming both MANU and MMUnlearner. This demonstrates that our method possesses a strong advantage in balancing unlearning effectiveness and model utility with respect to visual knowledge.
We present several examples where our approach effectively removes target visual knowledge while preserving non-target knowledge, as shown in Appendix. We compare our approach with GA, NPO, and MANU. Our method successfully erases the target visual content. In contrast, GA often leads to partial forgetting and distorted textual responses, whereas NPO tends to over-forget, degrading both visual and textual modalities.
4.2.2 Effect of VKD
The Visual Knowledge Distillation (VKD) module is designed to preserve the visual knowledge associated with non-target entities. Its strength is controlled by the parameter $\beta$ in Eq. 3. As shown in Fig. 2, when $\beta=0$—that is, when the VKD term in Eq. 3 is omitted—the visual knowledge related to non-target entities is not effectively preserved, as evidenced by the notably low accuracy on the Retain VQA task. In contrast, as $\beta$ increases, the accuracy on the Retain VQA set rises sharply, indicating a substantial improvement in preserving the visual knowledge associated with non-target entities.
Interestingly, our VKD module also contributes to forgetting the visual knowledge associated with the target entity, even though it was not originally designed for this purpose. As shown in Fig. 2, as $\beta$ increases, the accuracy on the Forget VQA set also exhibits a noticeable decline. These results illustrate that our approach effectively disentangles the forgetting of target visual knowledge from the preservation of non-target visual knowledge.
Notably, the strength of VKD should not be set excessively high. We observe that as $\beta$ increases, the accuracy on the Real-world VQA set initially rises but subsequently declines, indicating that an overly strong VKD may harm the model’s overall utility. Finally, we set the optimal strength to $\beta=0.3$, which achieves a balanced trade-off between the output-level objective and the VKD objective.
Moreover, based on the accuracy results on the Forget, Retain, and Real-world QA sets, we observe that they are only slightly influenced by the VKD module. This demonstrates that our VKD design does not compromise the preservation of textual knowledge.
4.2.3 Balance between Forgetting and Preservation
One inherent challenge in machine unlearning lies in balancing knowledge forgetting and preservation. Our approach employs two complementary strategies to address this issue. The first strategy involves selectively fine-tuning only a subset of parameters during unlearning, controlled by the mask m in Eq. 4. The second strategy introduces the weighting parameter $\alpha$ in Eq. 4 to regulate the relative strengths of forgetting and preservation objectives. We analyze the second strategy here and will discuss the first strategy in the following section.
The parameter $\alpha$ controls the strength of target visual knowledge forgetting relative to the preservation of other forms of knowledge, including the target textual knowledge (corresponding to $\mathcal{L}^{f}_{qa}$) and both the visual and textual knowledge associated with non-target entities (corresponding to $\mathcal{L}^{r}_{vqa}$ and $\mathcal{L}^{r}_{qa}$). As shown in Fig. 3, as $\alpha$ increases, the accuracy on the Forget VQA set drops rapidly, indicating that the forgetting of target visual knowledge is significantly enhanced. Meanwhile, the accuracies on the Retain VQA and Real-world VQA sets initially rise but subsequently decline, reflecting the inherent conflict between knowledge forgetting and preservation. Nevertheless, a proper forgetting strength of $\alpha=1.25$ achieves a desirable balance between the two objectives.
As shown in Fig. 3, the accuracies on the Forget, Retain, and Real-world QA sets remain largely insensitive to variations in $\alpha$, indicating that our approach effectively disentangles textual knowledge from visual knowledge.
Methods Fine-tuning Scope Running Time GPU memory
GA Full Model (Vision + LLM) 43.3 ($\pm$1.5)min 48.7 ($\pm$0.38)GB
GA_Diff Full Model (Vision + LLM) 57.4 ($\pm$1.5)min 50.1 ($\pm$0.37)GB
NPO Full Model (Vision + LLM) 61.3 ($\pm$1.5)min 68.3 ($\pm$0.38)GB
MANU Partial Layers 57.1 ($\pm$1.3)min 15.7 ($\pm$0.38)GB
MMU Full Model (Vision + LLM) 23.6 ($\pm$1.5)min 50.5 ($\pm$0.38)GB
Ours (VKD) Vision 16.7 ($\pm$1.3)min 47.7 ($\pm$0.24)GB
Ours (w/o VKD) Vision 15.4 ($\pm$1.3)min 46.7 ($\pm$0.24)GB
Table 2: Comparison of Unlearning Efficiency. Average running time per epoch (in minutes) on an NVIDIA RTX 4090 GPU. The GPU memory consumption is compared across these methods.
4.2.4 Efficiency of Unlearning
By freezing the large language model (LLM) and performing lightweight fine-tuning on the visual module, our method exhibits significant efficiency advantages. We measure the average per-epoch unlearning time under different configurations and compare it with other baseline methods on an NVIDIA 4090 GPU. As shown in Table. 2, our design achieves an excellent balance between forgetting precision and computational efficiency. Freezing the LLM reduces the overall unlearning time by more than 50%, while maintaining strong unlearning effectiveness. The introduction of VKD adds only minor overhead.
4.3 Ablation Study
Our approach incorporates several strategies to enhance MLLM unlearning, including partial fine-tuning, visual knowledge distillation, and the saliency masking scheme. We conduct an ablation study to demonstrate the contribution of each component and their complementary effects.
Unlike previous methods that fine-tune the entire MLLM or all MLP components during unlearning, we adopt a partial fine-tuning strategy: only the vision module—specifically, the MLP layers within the vision module—is fine-tuned, while the LLM backbone parameters remain frozen. We argue that this strategy offers efficiency advantages for unlearning, as the vision module is significantly smaller than the LLM backbone. Moreover, keeping the LLM frozen helps preserve the textual knowledge.
As shown in Table. 3, the baseline method (i.e., the “B” row) adopts a full fine-tuning strategy, wherein both the visual module and the LLM backbone are fine-tuned. When partial fine-tuning is excluded and only the VKD and masking schemes are applied (i.e., the “B+Mask+VKD” row), we observe that this configuration is inferior to our approach (i.e., the “F+Mask+VKD” row) in preserving textual knowledge, as evidenced by the lower accuracy on the Forget QA task. Moreover, its running time is significantly longer than that of our approach.
Comparing “F+Mask” to “F+VKD,” we observe that the masking scheme excels at preserving textual knowledge for both target and non-target entities, as reflected by the accuracy on the Forget and Retain QA sets. In contrast, “F+VKD” demonstrates superior performance in preserving visual knowledge associated with non-target entities and forgetting visual knowledge associated with the target entity. Thus, the masking and VKD strategies complement each other, enabling our approach (i.e., “F+Mask+VKD”) to outperform either strategy used individually.
We integrate two complementary schemes for selective forgetting: neuron pruning and weight masking. We conduct an ablation study by applying each scheme independently. The results demonstrate that our integration-based strategy outperforms any single scheme. Details of the ablation study are provided in Appendix.
Methods Forget Set Retain Set Realworld Set
Forget VQA Forget QA Retain VQA Retain QA RW VQA RW QA
Acc($\downarrow$) Acc($\uparrow$) Acc($\uparrow$) Acc($\uparrow$) Acc($\uparrow$) Acc($\uparrow$)
Vanilla 45.8 38.4 45.2 37.5 47.4 54.9
B 40.0 33.6 44.3 31.5 46.6 53.6
F 15.8 24.0 20.5 35.0 19.8 45.6
B+Mask 36.4 32.7 43.6 36.7 37.8 50.6
F+Mask 33.2 33.4 39.1 36.1 40.5 52.9
F+VKD 30.8 30.6 40.7 34.9 44.8 53.5
B+Mask+VKD 35.6 31.8 42.1 34.1 45.9 52.6
F+Mask+VKD 29.8 35.8 45.0 36.1 46.9 54.9
Table 3: Ablation study on the Forget/Retain/Realworld sets. Lower is better ($\downarrow$) for Forget VQA accuracy; higher is better ($\uparrow$) for the others. “B” denotes the baseline, which fine-tunes the entire MLLM model, whereas “F” denotes freezing the LLM while fine-tuning only the visual module. “F+mask+VKD” represents our proposed approach.
Task Method Original Epoch1 Epoch2 Epoch3 Epoch4 Epoch5 AG
Forget VQA MMU 29.1 31.2 35.4 37.4 36.2 37.4 8.3
MANU 38.6 38.2 40.8 44.6 42.9 42.9 4.3
Ours 29.3 29.2 27.6 27.6 30.6 30.6 1.3
Forget QA MMU 43.6 47.6 54.4 52.0 53.4 53.4 9.8
MANU 31.5 35.6 33.6 36.4 36.4 37.6 6.1
Ours 43.2 43.6 42.8 43.2 44.0 44.0 0.8
Table 4: Comparison of unlearning robustness for different methods over 5 epochs relearning attack.
4.4 Robustness of Unlearning
When conducting a relearning attack using a small portion of the forgetting data, we evaluate the extent to which the forgotten visual knowledge can be recovered after the attack. Specifically, we define a new metric, Accuracy Gap (AG), defined as the difference between the post-relearning accuracy and the accuracy of the original model, where the post-relearning accuracy refers to the final accuracy once relearning procedure has stably converged. The larger the AG, the more robust the unlearning algorithm is. In practice, we measure the AG using the accuracy on the Forget VQA task, i.e., by comparing the difference in Forget VQA accuracy before and after the relearning process.
We conduct relearning attacks using 10%, 20%, and 30% of the forgotten visual data. As shown in Table. 4, as the proportion of fine-tuning data used for the relearning attack increases, more forgotten knowledge is gradually recovered. However, with 20% of the relearning data, the Forget VQA accuracy shows only a slight recovery, increasing from 29.3% to 30.6% (i.e., AG=1.3%), demonstrating that our approach effectively forgets the undesired knowledge and remains robust against relearning attacks. More results are provided in Appendix.
5 Conclusion
In this paper, we propose an MLLM-oriented machine unlearning approach that disentangles visual and textual knowledge, handling them separately—specifically, erasing target visual knowledge while preserving textual knowledge. In addition to leveraging output-level supervision, we introduce the Visual Knowledge Distillation (VKD) scheme, which incorporates intermediate visual features from a teacher model as an auxiliary supervision signal. Interestingly, VKD not only facilitates forgetting of target visual knowledge but also aids in preserving visual knowledge of non-target entities. Moreover, we evaluate the robustness of our MLLM unlearning approach against relearning attacks, demonstrating that the forgotten visual knowledge cannot be easily recovered.

### v2

@@
-MLLM Machine Unlearning via Visual Knowledge Distillation
+Robust MLLM Unlearning via Visual Knowledge Distillation
 Abstract
-Recently, machine unlearning approaches have been proposed to remove sensitive information from well-trained large models.
-However, most existing methods are tailored for LLMs, while MLLM-oriented unlearning remains at its early stage.
-Inspired by recent studies exploring the internal mechanisms of MLLMs, we propose to disentangle the visual and textual knowledge embedded within MLLMs and introduce a dedicated approach to selectively erase target visual knowledge while preserving textual knowledge.
+Recently, LLM unlearning approaches have been proposed to remove sensitive information from well-trained large models.
+However, unlearning for Multimodal Large Language Models (MLLMs) remains at an early stage.
+Inspired by recent studies on the internal mechanisms of MLLMs, we propose to disentangle visual and textual knowledge within MLLMs and introduce a dedicated approach that selectively erases target visual knowledge while preserving textual knowledge.
 Unlike previous unlearning methods that rely on output-level supervision, our approach introduces a Visual Knowledge Distillation (VKD) scheme, which leverages intermediate visual representations within the MLLM as supervision signals.
 This design substantially enhances both unlearning effectiveness and model utility.
-Moreover, since our method only fine-tunes the visual components of the MLLM, it offers significant efficiency advantages.
+Moreover, since our method only fine-tunes the visual components of MLLM, it offers significant efficiency advantages.
 Extensive experiments demonstrate that our approach outperforms state-of-the-art unlearning methods in terms of both effectiveness and efficiency.
-Moreover, we are the first to evaluate the robustness of MLLM unlearning against relearning attacks.
+Furthermore, we are the first to evaluate the robustness of MLLM unlearning against re-learning attacks.
+Machine Learning, ICML
 1 Introduction
-Recently, Multimodal Large Language Models (MLLMs) have achieved remarkable progress, demonstrating outstanding performance on various multimodal applications such as image captioning and visual question answering (VQA) [liu2023visual, bai2023qwen].
+Recently, Multimodal Large Language Models (MLLMs) have achieved remarkable progress.
 However, growing concerns over data privacy have become a significant barrier to the widespread application of LLMs and MLLMs.
 This is because training large models typically involves the use of vast amounts of Internet data, which often contain sensitive personal information such as social security numbers or personal photographs.
-Moreover, numerous studies [pi2024mllm, li2024digger, cohen2024performance] have shown that large models can memorize portions of their training data and reproduce them verbatim in their outputs.
+Moreover, numerous studies (Pi et al., 2024; Li et al., 2024a; Cohen et al., 2024) have shown that large models can memorize portions of their training data and reproduce them verbatim in their outputs.
 This poses a substantial risk of privacy leakage.
-Therefore, the advent of data privacy and protection regulations—such as the European Union’s GDPR [regulation2018general] and the California Consumer Privacy Act (CCPA) [goldman2020introduction]—now mandates these large model providers to honor data deletion requests from individuals [dang2021right].
+Therefore, the advent of data privacy and protection regulations—such as the European Union’s GDPR (Regulation, General Data Protection, 2018) and the California Consumer Privacy Act (CCPA) (Goldman, 2020)—now mandates these large model providers to honor data deletion requests from individuals (Dang, 2021).
 The most straightforward approach to protecting privacy is to discard the trained model entirely, remove the individual’s data from the training set, and retrain a new model from scratch.
 However, retraining is computationally expensive and resource-intensive, particularly for large models.
-As a result, the field of machine unlearning [garg2020formalizing, ginart2019making, cao2015towards, gupta2021adaptive, sekhari2021remember, brophy2021machine, ullah2021machine] has emerged, focusing on the challenge of efficiently revising a trained model to forget specific data while preserving performance on the retaining dataset.
-Nowadays, many large-model-oriented machine unlearning methods have been proposed, such as Gradient Ascent (GA) [thudi2022unrollingsgd], its variations [liu2022privateunlearning], and Negative Preference Optimization (NPO) [zhang2024npo], etc.
-However, most existing approaches primarily focus on unlearning for LLMs, with relatively limited exploration in the context of MLLMs.
-In general, unlearning in LLMs is clearly defined—erasing the knowledge of a specific entity from the model.
-In contrast, defining unlearning for MLLMs is far more complex, as two distinct forms of knowledge are associated with the entity of interest: textual knowledge, stored within the LLM component, and visual knowledge, or visual patterns, linked to the entity’s appearance.
+As a result, the field of machine unlearning (Garg et al., 2020; Ginart et al., 2019; Cao and Yang, 2015; Gupta et al., 2021; Sekhari et al., 2021; Brophy and Lowd, 2021; Ullah et al., 2021) has emerged, aiming to efficiently revise trained models by selectively forgetting specific data while preserving performance on the remaining data.
+Nowadays, many LLM unlearning methods have been proposed, such as Gradient Ascent (GA) (Thudi et al., 2022), its variations (Liu et al., 2022), and Negative Preference Optimization (NPO) (Zhang et al., 2024), etc.
+However, MLLM unlearning remains at an early stage.
+LLM unlearning problem is clearly defined—erasing the knowledge of a specific entity from the model.
+In contrast, defining unlearning for MLLMs is far more complex, as two distinct forms of knowledge are associated with the entity of interest: textual knowledge, linked to the entity’s general information, and visual knowledge/visual patterns, linked to the entity’s appearance.
 For instance, for a well-known individual, the name (e.g., “Robin Williams”) and textual attributes (such as occupation or home address) constitute the textual knowledge, while the person’s facial appearance represents the visual knowledge.
-Consequently, there remains ambiguity in defining MLLM unlearning: some studies argue that it should involve erasing both textual and visual knowledge [liu2025MANU], whereas others contend that only the visual knowledge should be removed while preserving the textual knowledge [huo2025mmunlearner].
-In this paper, we follow [huo2025mmunlearner] in defining the desired MLLM unlearning objective as erasing visual knowledge while preserving textual knowledge, as this formulation offers greater flexibility.
+Consequently, there remains ambiguity in defining MLLM unlearning: some studies argue that it should involve erasing both textual and visual knowledge (Liu et al., 2025), whereas others contend that only the visual knowledge should be removed while preserving the textual knowledge (Huo et al., 2025).
+In this paper, we follow (Huo et al., 2025) to define the MLLM unlearning objective as erasing visual knowledge while preserving textual knowledge, as this formulation offers greater flexibility.
 The essence of this definition lies in disentangling visual and textual knowledge, thereby allowing selective removal of either component when necessary.
@@
 An MLLM typically consists of three components: a vision encoder, an LLM backbone, and a projector that bridges the two.
-Unlike [huo2025mmunlearner], which fine-tunes all three components for unlearning, we propose to fine-tune only the vision encoder and the projector while keeping the LLM backbone frozen.
-This design is motivated by some studies that explore the internal mechanisms of MLLMs [cohen2024performance, huang2024commonsenseknowledgeeditingbased, yu2024understanding].
+Unlike (Huo et al., 2025), which fine-tunes all three components for unlearning, we propose to fine-tune only the vision encoder and the projector while keeping the LLM backbone frozen.
+This design is motivated by some studies that explore the internal mechanisms of MLLMs (Cohen et al., 2024; Huang et al., 2024a; Yu and Ananiadou, 2024).
 These works decompose the VQA process into two stages: the first involves visually recognizing the entity (termed identification), and the second connects the recognized entity to its stored factual knowledge (termed extraction).
 It has been shown that, in a well-trained MLLM, the vision encoder and projector primarily handle the identification function, whereas the LLM module is responsible for the extraction function.
-Therefore, to effectively erase visual knowledge while preserving textual knowledge, we propose to keep the LLM module frozen during the unlearning process.
-Note that  [huo2025mmunlearner] conducted an ablation study comparing such partial fine-tuning setting (fine-tuning only the vision encoder and projector) with the full fine-tuning setting.
+Therefore, fine-tuning the visual module and disrupting the identification stage is sufficient to erase visual knowledge.
+Moreover, freezing the LLM component can largely preserve textual knowledge.
+Note that  (Huo et al., 2025) conducted an ablation study comparing such partial fine-tuning scheme (fine-tuning only the vision encoder and projector) with the full fine-tuning scheme.
 Their results demonstrated that the former is inferior to the latter in balancing unlearning effectiveness and model utility.
-In this paper, we propose a novel scheme, termed Visual Knowledge Distillation (VKD), to harness the advantages of our partial fine-tuning setting.
+In this paper, we propose a novel approach, termed Visual Knowledge Distillation (VKD), to harness the advantages of partial fine-tuning scheme.
 Specifically, when designing the unlearning objective function, most existing methods focus primarily on the model’s output.
 Take GA as an example: for the forgetting knowledge, it encourages the model to avoid producing the ground-truth answer, while for the remaining knowledge, it encourages the model to retain correct predictions.
-We argue that such output-level supervision is well-suited for full fine-tuning settings.
+We argue that such output-level supervision is well-suited for full fine-tuning scheme.
 However, in the partial fine-tuning scenario, this supervision signal must propagate through the frozen LLM, and thus becomes significantly weakened by the time it reaches the visual module.
-To address this issue, we propose to inject the supervision signal directly into the visual module.
+To address this issue, our VKD approach injects the supervision signal directly into the visual module.
 Concretely, we employ the original MLLM as the teacher model and use its intermediate visual representation (i.e., the output of the projector) as the supervision signal.
 We then introduce a loss term that encourages the unlearned model (the student model) to align with this signal.
-We term this approach Visual Knowledge Distillation (VKD), as it distills the teacher model’s visual representations—which predominantly encode visual knowledge—into the student model during unlearning.
+Our VKD distills the teacher model’s intermediate visual representations—which predominantly encode visual knowledge—into the student model during unlearning.
 The empirical results demonstrate that our approach outperforms previous methods in terms of both unlearning effectiveness and model utility.
@@
 Since the knowledge distillation is applied solely to the visual component, the fine-tuning process is highly efficient, requiring updates to only a small portion of parameters.
-Furthermore, due to its lightweight nature, our approach is well-suited for online unlearning scenarios, where data deletion requests occur frequently and arrive sequentially.
+Furthermore, due to its lightweight nature, our approach is well-suited for continual unlearning scenarios, where data deletion requests occur frequently and arrive sequentially.
 In such cases, our method can efficiently fine-tune the visual component on demand for each request.
 Beyond evaluating the effectiveness of unlearning, we also examine its robustness.
-Recent studies have revealed that machine unlearning is vulnerable to the so-called relearning attack, which attempts to recover forgotten data or knowledge using only a few samples from the forgetting set [lynch2024robustunlearning, hu2025joggingmemory].
+Recent studies have revealed that machine unlearning is vulnerable to the so-called relearning attack, which attempts to recover forgotten data or knowledge using only a few samples from the forgetting set (Lynch et al., 2024; Hu et al., 2025).
 If the ‘already-unlearned’ knowledge can be rapidly recovered, it suggests that the unlearning process merely obfuscates the knowledge rather than truly removes it.
@@
 Since LLMs may unintentionally memorize their pretraining data, Machine Unlearning (MU) techniques have been introduced to erase privacy-sensitive or copyrighted information from trained models.
-Gradient Ascent (GA) [thudi2022unrolling] is a representative method that fine-tunes the model to reduce the probability of generating sensitive information.
+Gradient Ascent (GA) (Thudi et al., 2022) is a representative method that fine-tunes the model to reduce the probability of generating sensitive information.
 One inherent challenge in unlearning lies in balancing unlearning effectiveness and model utility—that is, erasing the target data while preserving the remaining data.
-To address this, several GA variants have been proposed to mitigate the impact on remaining data by incorporating KL-divergence minimization [maini2024tofu] or by explicitly reducing the loss on the retained data [rafailov2023direct].
-Beyond supervised fine-tuning, several reinforcement learning-based methods, such as Negative Preference Optimization (NPO) [zhang2024negative], have been proposed to further enhance the effectiveness of unlearning.
+To address this, several GA variants have been proposed to mitigate the impact on remaining data by incorporating KL-divergence minimization (Maini et al., 2024) or by explicitly reducing the loss on the retained data (Rafailov et al., 2023).
+Beyond supervised fine-tuning, several reinforcement learning-based methods, such as Negative Preference Optimization (NPO) (Zhang et al., 2024), have been proposed to further enhance the effectiveness of unlearning.
 Machine Unlearning for MLLMs.
 Compared to MU for LLMs, research on MU for MLLMs remains in its infancy.
-SIU [li2024single] explores the erasure of visual patterns in MLLMs using real-world entity datasets through multifaceted fine-tuning.
-As noted by [yu2024understanding, cohen2024performance], MLLMs exhibit strong cross-modal interactions that entangle visual and textual knowledge, making selective unlearning more challenging and often resulting in substantial unintended knowledge loss.
-Moreover, even the definition of MLLM unlearning remains controversial: some studies define it as erasing visual knowledge while preserving textual knowledge [huo2025mmunlearner], whereas others argue it should involve forgetting both visual and textual knowledge [liu2025modality].
+SIU (Li et al., 2024b) explores the erasure of visual patterns in MLLMs using real-world entity datasets through multifaceted fine-tuning.
+As noted by (Yu and Ananiadou, 2024; Cohen et al., 2024), MLLMs exhibit strong cross-modal interactions that entangle visual and textual knowledge, making selective unlearning more challenging and often resulting in substantial unintended knowledge loss.
+Moreover, even the definition of MLLM unlearning remains controversial: some studies define it as erasing visual knowledge while preserving textual knowledge (Huo et al., 2025), whereas others argue it should involve forgetting both visual and textual knowledge (Liu et al., 2025).
 Nevertheless, existing methods predominantly rely on model outputs as the supervision signal for unlearning.
@@
 For the target entity, our objective is to (a) forget its visual knowledge while (b) preserving its textual knowledge.
-For non-target entities, we preserve both (c) the textual knowledge and (d) the visual knowledge through Visual Knowledge Distillation (VKD).
+For non-target entities, we preserve both (c) the textual knowledge and (d) the visual knowledge through VKD.
 3 Method
@@
 This design is motivated not only by efficiency considerations (as adjusting the visual module is significantly more lightweight than updating the entire MLLM) but also by the very definition of the MLLM unlearning objective.
-Since modern MLLMs are typically built upon a pretrained LLM backbone, it is widely assumed that textual knowledge is primarily encoded in the LLM, while visual knowledge is largely captured by the visual module [zhu2023minigpt, chen2024internvl].
+Since modern MLLMs are typically built upon a pretrained LLM backbone, it is widely assumed that textual knowledge is primarily encoded in the LLM, while visual knowledge is largely captured by the visual module (Zhu et al., 2023; Chen et al., 2024).
 Moreover, one of our core unlearning goals is to preserve textual knowledge while removing visual knowledge.
@@
 To this end, we introduce a new mechanism—Visual Knowledge Distillation (VKD)—to better balance unlearning and utility preservation.
-3.2.1 Visual Knowledge Distillation
+Visual Knowledge Distillation.
 It is important to note that all existing unlearning methods generally tackle this balance at the loss level: they maximize the loss on the forgetting dataset $D^{f}$ (to erase target knowledge) while minimizing the loss on the remaining dataset $D^{r}$ (to preserve other knowledge).
@@
 $\mathop{\min}_{\theta}(\mathcal{L}_{output}+\beta\mathcal{L}_{\text{VKD}})$ (3)
-[…]
+$\mathcal{L}_{\text{VKD}}=\|f^{r}_{vqa}-t^{r}_{vqa}\|^{2}$ (4)
 $\mathcal{L}_{output}=-\textbf{m}\odot\alpha\mathcal{L}^{f}_{vqa}+\mathcal{L}^{f}_{qa}+\mathcal{L}^{r}_{vqa}+\mathcal{L}^{r}_{qa}$ (5)
@@
 The parameter $\alpha$ is employed to trade off among the visual knowledge to be forgotten ($\mathcal{L}^{f}_{vqa}$), and the visual and textual knowledge to be preserved ($\mathcal{L}^{r}_{vqa}$, $\mathcal{L}^{f}_{qa}$ and $\mathcal{L}^{r}_{qa}$).
-3.2.2 Selective Forgetting
-As shown in Eq 4, we follow the method [huo2025mmunlearner] that leverage a masking scheme as performing unlearning, where m is a boolean mask that selectively updates parameters, and $\odot$ denotes the Hadamard product.
+Another advantage of intermediate-level supervision is that it eliminates the need to load the full reference model into memory; instead, only the visual module is required to obtain the supervision signal.
+Selective Forgetting.
+As shown in Eq 5, we follow the method (Huo et al., 2025) that leverage a masking scheme as performing unlearning, where m is a boolean mask that selectively updates parameters, and $\odot$ denotes the Hadamard product.
 Specifically, the mask m is referred to as the weight saliency map, as it quantifies how salient each parameter is with respect to a given dataset $D$.
-In addition to saliency-based masking, another selective forgetting technique is neuron pruning, which identifies neurons closely associated with the knowledge to be forgotten and achieves unlearning by pruning them.
-In this paper, we integrate these two selective forgetting strategies.
-They can be executed either in parallel or sequentially.
+In addition to saliency-based masking, another selective forgetting technique is neuron pruning (Pochinkov and Schoots, 2024), which identifies neurons associated with the knowledge to be forgotten and achieves unlearning by pruning them.
+In this paper, we integrate these two selective forgetting strategies into a unified framework.
+There are multiple ways to combine them, such as parallel or sequential integration.
 Our experiments demonstrate that the best performance is achieved when pruning is applied first, followed by masking-based fine-tuning.
-It is worth noting that in our approach, both neuron pruning and weight masking are applied exclusively to the visual module—specifically, the vision encoder $\mathcal{V}$ and the projector $\mathcal{P}$.
+Specifically, we adopt the neuron pruning strategy based on importance functions, as proposed in (Pochinkov and Schoots, 2024).
+We leverage the mean of absolute activation $I_{abs}$ as the importance function.
+Let $n$ be a neuron and denote its corresponding activations by $z$, the importance of neuron $n$ with respect to the dataset $D$ is defined as follows:
+$I_{\text{abs}}(D,n)=\frac{1}{|D|}\sum_{i\in D}|z(i)|$ (6)
+where $i\in D$ indicates the data sample $i$ in $D$.
+Thus, the neuron importance for unlearning $I_{u}$ is derived by the ratio of $I_{abs}$ between the knowledge to be forgotten and the knowledge to be retained:
+$I_{u}=\frac{I_{abs}(D^{f}_{vqa})}{I_{abs}(D^{f}_{qa})+I_{abs}(D^{r}_{vqa})+I_{abs}(D^{r}_{qa})}$ (7)
+Consequently, we identify the neurons to be pruned, denoted as $\textbf{n}_{P}$, by applying a threshold to $I_{u}$:
+$\textbf{n}_{P}=\mathbb{I}[I_{u}\geq d_{I}]$ (8)
+where $d_{I}$ is a predefined threshold.
+After selective neuron pruning, we perform masking-based fine-tuning.
+The key lies in computing the saliency of each model weight with respect to the knowledge to be forgotten.
+Previous studies (Huang et al., 2024b) have shown that this saliency can be approximated using the diagonal elements of the Fisher Information Matrix (FIM) of the vanilla model with respect to dataset $D$:
+$F^{D}_{diag}=[\Delta\mathcal{L}^{D}(\theta_{0})]^{2}$ (9)
+where $\theta_{0}$ denotes the vanilla model.
+Thus, the following ratio R indicates which parameters are salient for forgetting the target visual knowledge while preserving other knowledge:
+$R=\frac{[\Delta\mathcal{L}^{f}_{vqa}]^{2}}{[\Delta(\mathcal{L}^{f}_{qa}+\mathcal{L}^{r}_{vqa}+\mathcal{L}^{r}_{qa})]^{2}}$ (10)
+The Fisher mask $\textbf{m}_{F}$ is then obtained by comparing the ratio R against a predefined threshold $d_{F}$:
+$\textbf{m}_{F}=\mathbb{I}[R\geq d_{F}]$ (11)
+This mask is regarded as the final mask m used in Eq. 5.
+Besides this pruning-then-fine-tuning scheme, other variant strategies are discussed in Appendix.
+Unlearning on MLP layers.
+It is worth noting that in our approach, both neuron pruning and weight masking are applied to the visual module.
 This is motivated by recent studies on localizing factual memory within Transformer architectures.
-Prior research [liu2025MANU] has shown that MLP layers serve as primary knowledge storage components in Transformer architectures [geva2021transformerfeedforwardlayerskeyvalue, meng2023locatingeditingfactualassociations], whereas the self-attention modules mainly function as mechanisms for information propagation [elhage2021mathematical].
-In particular, for vision encoders such as the CLIP model, it has been observed that the high-dimensional inner projections within the feed-forward layers play a crucial role in storing visual knowledge, whereas the representations derived from the self-attention layers do not [ghiasi2022visiontransformerslearnvisual].
-Furthermore, it has been also observed that earlier layers learn textural attributes, whereas deeper layers learn high level object features or abstract concepts [ghiasi2022visiontransformerslearnvisual].
-Therefore, in our approach, both neuron pruning and weight masking are applied exclusively to the vision encoder’s MLP layers, with particular emphasis on the deeper layers.
-4 Evaluation
-4.1 Experimental Setup
-To accurately assess whether the target knowledge has been successfully unlearned from the MLLM, we typically do not perform unlearning directly on the original pretrained model (e.g., LLaVA-1.5-7B), as it is uncertain whether the model initially possesses the target knowledge.
-In practice, most unlearning benchmarks adopt a two-stage procedure: first, a fictitious dataset representing the target knowledge (e.g.,, synthetic personal profiles) is curated, and the original MLLM is fine-tuned on this dataset to obtain a vanilla model that explicitly contains the target knowledge.
-Then, the unlearning algorithm is applied to this vanilla model, and the resulting unlearned model is evaluated to assess the effectiveness of the unlearning process.
+Prior research (Liu et al., 2025) has shown that MLP layers serve as primary knowledge storage components in Transformer architectures (Geva et al., 2021; Meng et al., 2023), whereas the self-attention modules mainly function as mechanisms for information propagation (Elhage et al., 2021).
+In particular, for vision encoders such as the CLIP model, it has been observed that the high-dimensional inner projections within the feed-forward layers play a crucial role in storing visual knowledge, whereas the representations derived from the self-attention layers do not (Ghiasi et al., 2022).
+Furthermore, it has been observed that earlier layers learn textural attributes, whereas deeper layers learn high level object features or abstract concepts (Ghiasi et al., 2022).
+Therefore, in our approach, both neuron pruning and weight masking are applied to the vision encoder’s MLP layers, with emphasis on the deeper layers.
 Methods MLLMU-Bench CLEAR
@@
 GA_Diff 40.0 33.6 44.3 31.5 46.6 53.6 47.3 0.197 43.4 0.220 47.7 73.5
-KL_Min 42.4 33.6 44.9 32.0 47.4 54.6 40.4 0.270 38.1 0.274 51.5 82.8
-NPO 43.2 33.6 44.9 32.2 47.0 55.0 40.4 0.285 38.6 0.282 52.9 83.4
+KL_Min 42.4 33.6 44.9 32.0 46.6 54.6 40.4 0.270 38.1 0.274 51.5 82.8
+NPO 43.2 33.6 44.9 32.2 47.0 55.0 40.4 0.285 38.6 0.282 51.8 83.4
 MANU 41.6 33.6 43.2 32.1 46.3 54.8 39.8 0.263 25.8 0.224 49.8 68.3
-MMU 31.2 34.2 44.2 35.1 46.7 54.9 36.2 0.348 46.6 0.338 52.3 84.1
-Ours 29.8 35.8 45.0 36.1 46.9 54.9 32.7 0.327 49.6 0.346 52.3 84.8
+MMU 31.2 34.2 44.2 35.1 46.7 54.9 36.2 0.348 46.6 0.338 51.2 84.1
+Ours 29.8 35.8 45.0 36.1 47.2 54.9 34.2 0.349 49.6 0.346 52.8 84.8
 Qwen2-VL-7B
@@
 GA 50.4 46.7 51.5 57.6 74.4 77.8 55.3 0.123 62.4 0.083 65.9 86.8
-GA_Diff 54.4 52.8 38.8 54.4 74.5 77.0 63.3 0.125 71.4 0.088 69.0 92.7
-KL_Min 45.6 45.3 35.9 55.6 74.8 77.1 67.0 0.120 70.9 0.098 68.4 90.7
+GA_Diff 54.4 52.8 38.8 54.4 74.5 77.0 63.3 0.125 64.4 0.088 66.0 92.7
+KL_Min 45.6 45.3 35.9 55.6 74.8 77.1 67.0 0.120 68.9 0.098 68.4 90.7
 NPO 49.6 50.4 49.5 53.3 75.2 78.3 62.8 0.103 68.3 0.091 68.9 88.7
 MANU 43.2 47.2 43.2 49.4 73.7 76.2 61.3 0.107 65.2 0.094 66.2 87.3
-MMU 44.0 54.4 56.0 55.7 75.3 77.3 50.0 0.123 70.9 0.100 68.9 94.7
-Ours 42.8 54.6 56.0 56.4 74.9 77.4 48.7 0.125 71.5 0.094 69.0 90.8
+MMU 44.0 54.4 52.6 55.7 75.3 77.3 50.0 0.123 68.3 0.100 68.9 94.7
+Ours 42.8 54.6 56.0 56.4 75.4 77.4 48.7 0.125 70.9 0.094 69.0 90.8
 Table 1: Overall results of baselines and our approach on MLLMU-Bench and CLEAR benchmarks.
-$\downarrow$ Lower is better for Forget VQA accuracy (reflecting better forgetting of target visual knowledge), $\uparrow$ Higher is better for the others (reflecting better retention of non-target visual knowledge and textual knowledge).
+Bold and underlined values indicate the best and second-best performance, respectively.
+$\downarrow$ Lower is better for Forget VQA accuracy, $\uparrow$ Higher is better for the others.
 For MLLMU-Bench, the Forget/Retain QA tasks are evaluated using the Acc metric, while for CLEAR, these tasks are evaluated using the ROUGE-L metric.
-4.1.1 Datasets and Metrics
+4 Evaluation
+To accurately assess whether the target knowledge has been successfully unlearned from the MLLM, we typically do not perform unlearning directly on the original pretrained model (e.g., LLaVA-1.5-7B), as it is uncertain whether the model initially possesses the target knowledge.
+In practice, most unlearning benchmarks adopt a two-stage procedure: first, a fictitious dataset representing the target knowledge (e.g., synthetic personal profiles) is curated, and the base MLLM is fine-tuned on this dataset to obtain a vanilla model that explicitly contains the target knowledge.
+Then, the unlearning algorithm is applied to this vanilla model, and the resulting unlearned model is evaluated to assess the effectiveness of the unlearning algorithm.
+4.1 Datasets and Metrics
 Datasets: We employ two benchmark datasets to evaluate unlearning performance in MLLMs.
@@
 CLEAR is built on top of an LLM unlearning dataset, i.e., the TOFU dataset.
-For each author in TOFU, CLEAR adds several face images and corresponding textual descriptions generated by GPT-4o [achiam2023gpt].
+For each author in TOFU, CLEAR adds several face images and corresponding textual descriptions generated by GPT-4o (Achiam et al., 2023).
 Similar to MLLMU-Bench, CLEAR also comprises Forget, Retain, and Real-world sets.
 Metrics: We evaluate the forgetting and retention of textual knowledge using the Question Answering (QA) task, and assess the forgetting and retention of visual knowledge using the Visual Question Answering (VQA) task.
-The performance on multiple-choice questions is measured by average accuracy, while text generation performance is evaluated using the ROUGE-L metric [lin2004rouge].
+The performance on multiple-choice questions is measured by average accuracy, while text generation performance is evaluated using the ROUGE-L metric (Lin, 2004).
 Beyond evaluating the effectiveness of machine unlearning, we further investigate its robustness.
@@
 Specifically, after performing MLLM unlearning, we fine-tune the unlearned model with a small subset of the forgotten visual knowledge and subsequently assess the extent to which this forgotten knowledge can be recovered through a relearning process.
-4.1.2 Models and Baselines
+4.2 Models and Baselines
 MLLMs: We conduct experiments using open-source MLLMs as base models.
@@
 The vanilla models are trained following the official implementations provided by MLLMU-Bench and CLEAR.
-For each vanilla model, we perform selective unlearning of visual knowledge according to our proposed framework, while freezing the LLM weights.
-Baselines: We compare our method with six baseline approaches: GA [thudi2022unrolling], which applies gradient ascent on the forget VQA set $D_{f}$; GA_Diff [liu2022continual], an improved GA variant with a joint loss to balance the forget set $D_{f}$ and retain set $D_{r}$; KL_Min [maini2024tofu], which aligns predictions on $D_{r}$ with the original model while maximizing KL divergence with $D_{f}$; NPO [zhang2024negative], which treats $D_{f}$ as negative-preference data and integrates unlearning into a preference optimization framework with an oracle model fine-tuned on $D_{r}$; MANU [liu2025modality], which introduces modality-aware neuron unlearning by selectively pruning neurons that contribute most to the forget set, preserving model performance while achieving balanced forgetting across modalities; MMUnlearner [huo2025mmunlearner], which uses saliency-weighted updates to selectively protect non-target parameters while forgetting visual knowledge.
-All implementations are based on official MLLMU-Bench and CLEAR code, using the same learning rate, forgetting rounds, batch size, and full-precision training on four NVIDIA 4090 GPUs.
+Baselines: We compare our method with six baseline approaches: GA (Thudi et al., 2022), which applies gradient ascent on the forget VQA set $D_{f}$; GA_Diff (Liu et al., 2022), an improved GA variant with a joint loss to balance the forget set $D_{f}$ and retain set $D_{r}$; KL_Min (Maini et al., 2024), which aligns predictions on $D_{r}$ with the original model while maximizing KL divergence with $D_{f}$; NPO (Zhang et al., 2024), which treats $D_{f}$ as negative-preference data and integrates unlearning into a preference optimization framework with an oracle model fine-tuned on $D_{r}$; MANU (Liu et al., 2025), which introduces modality-aware neuron unlearning by selectively pruning neurons that contribute most to the forget set, preserving model performance while achieving balanced forgetting across modalities; MMU(nlearner) (Huo et al., 2025), which uses saliency-weighted updates to selectively protect non-target parameters while forgetting visual knowledge.
 Figure 2: The trade-off between the output-level objective and the VKD objective with respect to the VKD strength $\beta$.
-Figure 3: The trade-off between knowledge forgetting and preservation with respect to the weighting parameter $\alpha$.
 4.3 Main Results and Analysis
-4.2.1 Effectiveness of Unlearning
-On MLLMU-Bench, our method achieves the lowest accuracy on the Forget VQA task for both LLaVA-7B and Qwen2-VL, surpassing the state-of-the-art method MMUnlearner by 1.4% and 1.2%, respectively.
-Similarly, on CLEAR, our approach consistently surpasses MMUnlearner by 3.5% and 1.3%, respectively.
+Effectiveness of Unlearning.
+As shown in Table 1, our approach can significantly reduce the accuracy of Forget VQA task on MLLMU-Bench for both LLaVA-7B and Qwen2-VL, surpassing the state-of-the-art method MMUnlearner by 1.4% and 1.2%, respectively.
+Similarly, on CLEAR, our approach consistently surpasses MMUnlearner by 2.0% and 1.3%, respectively.
 This illustrates that our approach can effectively erase the visual knowledge associated with the target entity.
-Our method also achieves state-of-the-art performance on both the Retain QA task and the Forget QA task, particularly in MLLMU-Bench dataset.
+Our method also retains the accuracy of both the Retain QA task and the Forget QA task, particularly in MLLMU-Bench dataset.
 This demonstrates that our approach effectively preserves the textual knowledge associated with both target and non-target entities.
-Compared to textual knowledge forgetting, retaining the visual knowledge associated with non-target entities remains the most challenging issue.
+Compared to target visual knowledge forgetting, retaining the visual knowledge associated with non-target entities remains the most challenging issue.
 From the results on the Retain VQA and Real-world VQA, our approach shows substantial improvement, outperforming both MANU and MMUnlearner.
 This demonstrates that our method possesses a strong advantage in balancing unlearning effectiveness and model utility with respect to visual knowledge.
-We present several examples where our approach effectively removes target visual knowledge while preserving non-target knowledge, as shown in Appendix.
-We compare our approach with GA, NPO, and MANU.
-Our method successfully erases the target visual content.
+We present several examples for comparing our approach with GA, NPO, and MANU, as shown in Appendix.
+We find that our method can successfully erases the target visual content.
 In contrast, GA often leads to partial forgetting and distorted textual responses, whereas NPO tends to over-forget, degrading both visual and textual modalities.
-4.2.2 Effect of VKD
-The Visual Knowledge Distillation (VKD) module is designed to preserve the visual knowledge associated with non-target entities.
+Figure 3: The trade-off between knowledge forgetting and preservation with respect to the weighting parameter $\alpha$.
+Effect of VKD.
+The VKD module is designed to preserve the visual knowledge associated with non-target entities.
 Its strength is controlled by the parameter $\beta$ in Eq.
@@
 3 is omitted—the visual knowledge related to non-target entities is not effectively preserved, as evidenced by the notably low accuracy on the Retain VQA task.
-In contrast, as $\beta$ increases, the accuracy on the Retain VQA set rises sharply, indicating a substantial improvement in preserving the visual knowledge associated with non-target entities.
+In contrast, as $\beta$ increases, the accuracy of the Retain VQA rises sharply, indicating a substantial improvement in preserving the visual knowledge associated with non-target entities.
 Interestingly, our VKD module also contributes to forgetting the visual knowledge associated with the target entity, even though it was not originally designed for this purpose.
 As shown in Fig.
-2, as $\beta$ increases, the accuracy on the Forget VQA set also exhibits a noticeable decline.
+2, as $\beta$ increases, the accuracy of the Forget VQA also exhibits a noticeable decline.
 These results illustrate that our approach effectively disentangles the forgetting of target visual knowledge from the preservation of non-target visual knowledge.
 Notably, the strength of VKD should not be set excessively high.
-We observe that as $\beta$ increases, the accuracy on the Real-world VQA set initially rises but subsequently declines, indicating that an overly strong VKD may harm the model’s overall utility.
+We observe that as $\beta$ increases, the accuracy of the Real-world VQA initially rises but subsequently declines, indicating that an overly strong VKD may harm the model’s overall utility.
 Finally, we set the optimal strength to $\beta=0.3$, which achieves a balanced trade-off between the output-level objective and the VKD objective.
-Moreover, based on the accuracy results on the Forget, Retain, and Real-world QA sets, we observe that they are only slightly influenced by the VKD module.
+Moreover, based on the accuracy of the Forget QA, Retain QA, and Real-world QA, we observe that they are only slightly influenced by the VKD module.
 This demonstrates that our VKD design does not compromise the preservation of textual knowledge.
-4.2.3 Balance between Forgetting and Preservation
+Balance between Forgetting and Preservation.
 One inherent challenge in machine unlearning lies in balancing knowledge forgetting and preservation.
@@
 The first strategy involves selectively fine-tuning only a subset of parameters during unlearning, controlled by the mask m in Eq.
-4.
+5.
 The second strategy introduces the weighting parameter $\alpha$ in Eq.
@@
 We analyze the second strategy here and will discuss the first strategy in the following section.
-The parameter $\alpha$ controls the strength of target visual knowledge forgetting relative to the preservation of other forms of knowledge, including the target textual knowledge (corresponding to $\mathcal{L}^{f}_{qa}$) and both the visual and textual knowledge associated with non-target entities (corresponding to $\mathcal{L}^{r}_{vqa}$ and $\mathcal{L}^{r}_{qa}$).
-As shown in Fig.
-3, as $\alpha$ increases, the accuracy on the Forget VQA set drops rapidly, indicating that the forgetting of target visual knowledge is significantly enhanced.
-Meanwhile, the accuracies on the Retain VQA and Real-world VQA sets initially rise but subsequently decline, reflecting the inherent conflict between knowledge forgetting and preservation.
-Nevertheless, a proper forgetting strength of $\alpha=1.25$ achieves a desirable balance between the two objectives.
-As shown in Fig.
-3, the accuracies on the Forget, Retain, and Real-world QA sets remain largely insensitive to variations in $\alpha$, indicating that our approach effectively disentangles textual knowledge from visual knowledge.
 Methods Fine-tuning Scope Running Time GPU memory
@@
 NPO Full Model (Vision + LLM) 61.3 ($\pm$1.5)min 68.3 ($\pm$0.38)GB
-MANU Partial Layers 57.1 ($\pm$1.3)min 15.7 ($\pm$0.38)GB
 MMU Full Model (Vision + LLM) 23.6 ($\pm$1.5)min 50.5 ($\pm$0.38)GB
@@
 Ours (w/o VKD) Vision 15.4 ($\pm$1.3)min 46.7 ($\pm$0.24)GB
-Table 2: Comparison of Unlearning Efficiency. Average running time per epoch (in minutes) on an NVIDIA RTX 4090 GPU. The GPU memory consumption is compared across these methods.
-4.2.4 Efficiency of Unlearning
-By freezing the large language model (LLM) and performing lightweight fine-tuning on the visual module, our method exhibits significant efficiency advantages.
-We measure the average per-epoch unlearning time under different configurations and compare it with other baseline methods on an NVIDIA 4090 GPU.
-As shown in Table.
-2, our design achieves an excellent balance between forgetting precision and computational efficiency.
+Table 2: Comparison of Unlearning Efficiency. Average running time per epoch (in minutes) on NVIDIA RTX 4090 GPU. The GPU memory consumption is compared across these methods.
+The parameter $\alpha$ controls the strength of target visual knowledge forgetting relative to the preservation of other forms of knowledge, including the target textual knowledge (corresponding to $\mathcal{L}^{f}_{qa}$) and both the visual and textual knowledge associated with non-target entities (corresponding to $\mathcal{L}^{r}_{vqa}$ and $\mathcal{L}^{r}_{qa}$).
+As shown in Fig.
+3, as $\alpha$ increases, the accuracy of the Forget VQA drops rapidly, indicating that the forgetting of target visual knowledge is significantly enhanced.
+Meanwhile, the accuracies of the Retain VQA and Real-world VQA initially rise but subsequently decline, reflecting the inherent conflict between knowledge forgetting and preservation.
+Nevertheless, a proper forgetting strength of $\alpha=1.25$ achieves a desirable balance between the two objectives.
+As shown in Fig.
+3, the accuracies of the Forget QA, Retain QA, and Real-world QA remain largely insensitive to variations in $\alpha$, indicating that our approach effectively disentangles textual knowledge from visual knowledge.
+Efficiency of Unlearning.
+By freezing the LLM backbone and applying lightweight fine-tuning to the visual module, our method achieves notable efficiency gains.
+We measure the average per-epoch unlearning time under different configurations and compare with other baseline methods on NVIDIA 4090 GPU.
+As shown in Table 2, our approach achieves a favorable trade-off between unlearning efficacy and computational efficiency.
 Freezing the LLM reduces the overall unlearning time by more than 50%, while maintaining strong unlearning effectiveness.
 The introduction of VKD adds only minor overhead.
-4.3 Ablation Study
-Our approach incorporates several strategies to enhance MLLM unlearning, including partial fine-tuning, visual knowledge distillation, and the saliency masking scheme.
-We conduct an ablation study to demonstrate the contribution of each component and their complementary effects.
-Unlike previous methods that fine-tune the entire MLLM or all MLP components during unlearning, we adopt a partial fine-tuning strategy: only the vision module—specifically, the MLP layers within the vision module—is fine-tuned, while the LLM backbone parameters remain frozen.
-We argue that this strategy offers efficiency advantages for unlearning, as the vision module is significantly smaller than the LLM backbone.
-Moreover, keeping the LLM frozen helps preserve the textual knowledge.
-As shown in Table.
-3, the baseline method (i.e., the “B” row) adopts a full fine-tuning strategy, wherein both the visual module and the LLM backbone are fine-tuned.
-When partial fine-tuning is excluded and only the VKD and masking schemes are applied (i.e., the “B+Mask+VKD” row), we observe that this configuration is inferior to our approach (i.e., the “F+Mask+VKD” row) in preserving textual knowledge, as evidenced by the lower accuracy on the Forget QA task.
-Moreover, its running time is significantly longer than that of our approach.
-Comparing “F+Mask” to “F+VKD,” we observe that the masking scheme excels at preserving textual knowledge for both target and non-target entities, as reflected by the accuracy on the Forget and Retain QA sets.
-In contrast, “F+VKD” demonstrates superior performance in preserving visual knowledge associated with non-target entities and forgetting visual knowledge associated with the target entity.
-Thus, the masking and VKD strategies complement each other, enabling our approach (i.e., “F+Mask+VKD”) to outperform either strategy used individually.
-We integrate two complementary schemes for selective forgetting: neuron pruning and weight masking.
-We conduct an ablation study by applying each scheme independently.
-The results demonstrate that our integration-based strategy outperforms any single scheme.
-Details of the ablation study are provided in Appendix.
 Methods Forget Set Retain Set Realworld Set
@@
 F+Mask+VKD 29.8 35.8 45.0 36.1 46.9 54.9
-Table 3: Ablation study on the Forget/Retain/Realworld sets.
+Table 3: Ablation study.
 Lower is better ($\downarrow$) for Forget VQA accuracy; higher is better ($\uparrow$) for the others.
@@
 Table 4: Comparison of unlearning robustness for different methods over 5 epochs relearning attack.
+4.4 Ablation Study
+Our approach incorporates several strategies to enhance MLLM unlearning, including partial fine-tuning, visual knowledge distillation, and the saliency masking scheme.
+We conduct an ablation study to demonstrate the contribution of each component and their complementary effects.
+Unlike previous methods that fine-tune the entire MLLM or all MLP components during unlearning, we adopt a partial fine-tuning strategy: only the vision module—specifically, the MLP layers within the vision module—is fine-tuned, while the LLM backbone parameters remain frozen.
+We argue that this strategy offers efficiency advantages for unlearning, as the vision module is significantly smaller than the LLM backbone.
+Moreover, keeping the LLM frozen helps preserve the textual knowledge.
+As shown in Table 3, the baseline method (i.e., the “B” row) adopts a full fine-tuning strategy, wherein both the visual module and the LLM backbone are fine-tuned.
+When partial fine-tuning is excluded and only the VKD and masking schemes are applied (i.e., the “B+Mask+VKD” row), we observe that this configuration is inferior to our approach (i.e., the “F+Mask+VKD” row) in preserving textual knowledge, as evidenced by the lower accuracy on the Forget QA task.
+Moreover, its running time is significantly longer than that of our approach.
+Comparing “F+Mask” to “F+VKD,” we observe that the masking scheme excels at preserving textual knowledge for both target and non-target entities, as reflected by the accuracy on the Forget and Retain QA sets.
+In contrast, “F+VKD” demonstrates superior performance in preserving visual knowledge associated with non-target entities and forgetting visual knowledge associated with the target entity.
+Thus, the masking and VKD strategies complement each other, enabling our approach (i.e., “F+Mask+VKD”) to outperform either strategy used individually.
+We integrate two complementary schemes for selective forgetting: neuron pruning and weight masking.
+We conduct an ablation study by applying each scheme independently.
+The results demonstrate that our integration-based strategy outperforms any single scheme.
+Details of the ablation study are provided in Appendix.
 4.5 Robustness of Unlearning
@@
 Specifically, we define a new metric, Accuracy Gap (AG), defined as the difference between the post-relearning accuracy and the accuracy of the original model, where the post-relearning accuracy refers to the final accuracy once relearning procedure has stably converged.
-The larger the AG, the more robust the unlearning algorithm is.
+The smaller the AG, the more robust the unlearning algorithm is.
 In practice, we measure the AG using the accuracy on the Forget VQA task, i.e., by comparing the difference in Forget VQA accuracy before and after the relearning process.
 We conduct relearning attacks using 10%, 20%, and 30% of the forgotten visual data.
-As shown in Table.
-4, as the proportion of fine-tuning data used for the relearning attack increases, more forgotten knowledge is gradually recovered.
+As shown in Table 4, as the proportion of fine-tuning data used for the relearning attack increases, more forgotten knowledge is gradually recovered.
 However, with 20% of the relearning data, the Forget VQA accuracy shows only a slight recovery, increasing from 29.3% to 30.6% (i.e., AG=1.3%), demonstrating that our approach effectively forgets the undesired knowledge and remains robust against relearning attacks.
-More results are provided in Appendix.
+More results are provided in Appendix C.
 5 Conclusion
-In this paper, we propose an MLLM-oriented machine unlearning approach that disentangles visual and textual knowledge, handling them separately—specifically, erasing target visual knowledge while preserving textual knowledge.
-In addition to leveraging output-level supervision, we introduce the Visual Knowledge Distillation (VKD) scheme, which incorporates intermediate visual features from a teacher model as an auxiliary supervision signal.
+In this paper, we propose an MLLM unlearning approach that fine-tunes only the visual module while keeping the LLM backbone frozen.
+We introduce a VKD scheme, which incorporates intermediate visual representation from a teacher model as a supervision signal.
 Interestingly, VKD not only facilitates forgetting of target visual knowledge but also aids in preserving visual knowledge of non-target entities.
+Meanwhile, we combine neuron pruning and FIM masking in a unified framework.
 Moreover, we evaluate the robustness of our MLLM unlearning approach against relearning attacks, demonstrating that the forgotten visual knowledge cannot be easily recovered.